OpenStudy (anonymous):

Find the slope and y intercept for the given line. 3 x + 2 y = 7 m = y -intercept Find the slope and y intercept for the given line. 3 x + 2 y = 7 m = y -intercept @Mathematics

OpenStudy (anonymous):

3x + 2y = 7 2y = 7 - 3x y = (7 - 3x)/2 y = 3.5 - 1.5x intercept = 3.5 slope = -1.5